# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 2.1 Mathematical notation
Throughout the article, Hilbert spaces are denoted by $\mathcal{H}$, and are assumed to be finite-dimensional unless explicitly stated otherwise. Elements of $\mathcal{H}$ are denoted by $|v\rangle\in\mathcal{H}$, where the inner product $\langle u|v\rangle$ for $|v\rangle,|u\rangle\in\mathcal{H}$ is linear in the second argument and defines the vector norm $\||v\rangle\|=\sqrt{\langle v|v\rangle}$. Quantum states are the norm 1 elements of $\mathcal{H}$. In this work, $\mathbb{B}(\mathcal{H})$ denotes the unital $\dagger$-algebra of bounded linear operators on $\mathcal{H}$ with norm $\|M\|^{2}_{\mathrm{op}}=\sup_{|v\rangle\in\mathcal{H}}\langle v|M^{\dagger}M|v\rangle/\langle v|v\rangle$. The unit in $\mathbb{B}(\mathcal{H})$ is denoted by $\mathbbm{1}$, and we write $|M|=M^{\dagger}M$ for the positive part of $M\in\mathbb{B}(\mathcal{H})$. Given a set $A$, a collection of positive operators $\{M_{a}\geq 0:a\in\mathcal{A}\}$ with the property that $\sum_{a}M_{a}=\mathbbm{1}$, is called a POVM over $\mathcal{A}$. When the operators in a POVM are orthogonal projections, we call it a PVM. Given another finite set $\mathcal{X}$, we let $\mathscr{A}_{POVM}^{(\mathcal{X},\mathcal{A})}$ be the $\dagger$-subalgebra of $\mathbb{B}(\mathcal{H})$ consisting of $\dagger$-polynomials in the POVM elements $\{M_{a|x}\geq 0:a\in\mathcal{A}\}$. Given a random variable $X$ with a distribution $\mu:X\to\mathbb{R}_{\geq 0}$ such that $\sum_{x\in X}\mu(x)=1$ we denote the expectation of $X$ by $\mathop{\mathbb{E}}[X]=\sum_{x\in X}\mu(x)\cdot x$. For $a,b\in\mathbb{R}$ and $\delta>0$, $a\approx_{\delta}b$ is short for $|a-b|\leq\delta$.
2.2 Efficient quantum circuits and homomorphic encryption
To define the KLVY compilation procedure we require the following concepts from quantum cryptography.
Definition 2.1.
A procedure $\mathcal{P}$ is quantum polynomial time (QPT) if:
  1. 1.
  2. 2.
In particular, the following definition is relevant in Section 3.
Definition 2.2.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be Hilbert spaces. An isometry $V:\mathcal{H}\mapsto\mathcal{H}^{\prime}$ is efficient if there is a QPT procedure implementing $V$.
For more details on efficiently implementable quantum operations we refer the reader to the discussions in [CMM+24]. We also need to define a quantum homomorphic encryption (QHE) scheme. A formal definition of QHE first appeared in [BJ15]. We follow the description of QHE outlined in [KLVY23, Bra18]:
Definition 2.3.
  1. (i)
  2. (ii)
  3. (iii)
    A QPT algorithm $\mathsf{Eval}$ which takes as input a classical description of a quantum circuit $\mathsf{C}:\mathcal{H}\otimes(\mathbb{C}^{2})^{\otimes n}\to(\mathbb{C}^{2})^{\otimes m}$ from $\mathcal{C}$, a quantum plaintext $|\Psi\rangle\in\mathcal{H}$ on Hilbert space, a ciphertext $\chi$, and evaluates a quantum circuit $\mathsf{Eval}_{\mathsf{C}}(|\Psi\rangle\otimes|0\rangle^{\mathrm{poly}(\lambda,n)},\chi)$ producing a ciphertext $\alpha\in\{0,1\}^{\ell}$.
  4. (iv)
  1. 1.
    1. (a)
    2. (b)
  2. 2.
    (IND-CPA security). Fix a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$. Any quantum polynomial time adversary $\mathfrak{A}$ with access to $\mathsf{Enc}(\cdot,\mathsf{sk})$ (but does not know $\mathsf{sk}$) cannot distinguish between ciphertexts $\chi\leftarrow\mathsf{Enc}(x_{0},\mathsf{sk})$ and $\chi^{\prime}\leftarrow\mathsf{Enc}(x_{1},\mathsf{sk})$ with non-negligible probability in $\lambda$, where $x_{0}$ and $x_{1}$ are any plaintexts chosen by the adversary. That is
    for all pairs $(x_{0},x_{1})$.
  3. 3.
    (Compactness). The runtime of $\mathsf{Dec}$ does not depend on the description of $\mathsf{C}$. That is, $|\mathsf{Eval}_{C}(|\Psi\rangle\otimes|0\rangle^{\mathrm{poly}(\lambda,n)},\chi)|=\mathrm{poly}(\lambda)$.
  4. 4.
    (Universality). There is a QHE scheme with (i)-(iv) for any quantum circuit.
For the KLVY compilation procedure, we only require schemes that satisfy (1) and (2). A QHE satisfying (3) and (4) is called a quantum fully homomorphic encryption (QFHE) scheme. Schemes satisfying (1)-(4) have been proposed in [Mah20, Bra18].
@@ 2.3 Bell scenarios, inequalities, and violations
Let $\mathcal{A},\mathcal{B},\mathcal{X},$ and $\mathcal{Y}$ be finite sets, with $|\mathcal{A}|=m_{A}$, $|\mathcal{B}|=m_{B}$, $|\mathcal{X}|=n_{A}$, and $|\mathcal{Y}|=n_{B}$. A bipartite Bell scenario is described by the tuple $\mathcal{S}=(\mathcal{A},\mathcal{B},\mathcal{X},\mathcal{Y},\pi)$, where $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a distribution over the measurement settings. In this scenario, each party receives an input $x\in\mathcal{X}$ (resp. $y\in\mathcal{Y}$) sampled according to $\pi$, and returns outputs $a\in\mathcal{A}$ (resp. $b\in\mathcal{B}$). The parties are non-communicating, and therefore cannot coordinate their outputs. The behaviour of the provers is characterized by a correlation, a set of conditional probabilities $\mathsf{p}=\{p(a,b|x,y):a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$, which is realized by an underlying physical theory or model. In the quantum setting, we allow the provers to share a bipartite quantum state, and say the correlation $\mathsf{p}$ is realized by a bipartite (quantum) model
@@
where $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ are Hilbert spaces, $\{M_{a|x}\}_{a\in\mathcal{A}}$ and $\{N_{b|y}\}_{b\in\mathcal{B}}$ are POVMs on $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ respectively, and $|\Psi\rangle_{AB}$ is a vector state in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. More generally, a correlation $\mathsf{p}$ is quantum (or an element of $C_{\mathrm{q}}(n_{A},n_{B},m_{A},m_{B})$) if there exists a bipartite model $\mathrm{Q}$ for which $\mathsf{p}$ can be realized via the Born rule as
$p(a,b,|x,y)=\langle\Psi|M_{a|x}\otimes N_{b|y}|\Psi\rangle.$ (2.2)
We denote the class of bipartite (quantum) models by $\mathcal{Q}(n_{A},n_{B},m_{A},m_{B})$. From now on we will refer to such models simply as bipartite models.
In contrast to the set of quantum correlations, we have the collection of local correlations $C_{\mathrm{loc}}(n_{A},n_{B},m_{A},m_{B})$. These are the correlations $\{p(a,b,|x,y)\}$ for which there exists a classical model, that is a probability distribution $\lambda_{k}$ and a local distributions $p_{k}^{A}(a|x)$ and $p_{k}^{B}(b|y)$ such that
$p(a,b,|x,y)=\sum_{k}\lambda_{k}\,p_{k}^{A}(a|x)\,p_{k}^{B}(b|y).$ (2.3)
We let $\mathrm{C}=(\lambda_{k},\{p_{k}^{A}\},\{p_{k}^{B}\})$ denote a classical model and let $\mathcal{L}(n_{A},n_{B},m_{A},m_{B})$ denote the class of all classical models. In what follows we consider Bell scenarios where $n_{A}=n_{B}=n$, and $m_{A}=m_{B}=m$. Bell’s theorem [Bel64] states that $C_{loc}(2,2)$ is a strict subset of $C_{q}(2,2)$.
@@
$I=\sum_{a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}}w_{abxy}\,p(a,b,|x,y),$ (2.4)
@@
for coefficients $w_{abxy}\in\mathbb{R}$. A Bell inequality is a functional $I$ and a bound $\eta>0$ such that $I\leq\eta$ for all $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. Given a functional $I$, the classical value is the maximal value achieved by the classical correlations $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. We denote this value by $\eta^{\mathrm{L}}\coloneqq\sup_{\mathsf{p}\in C_{\mathrm{loc}}(m,n)}I$. The quantum value for $I$ is the maximal value achieved by the set of quantum correlations $\mathsf{p}\in C_{\mathrm{q}}(m,n)$, and we denote the quantum value on $I$ by $\eta^{\mathrm{Q}}\coloneqq\sup_{\mathsf{p}\in C_{\mathrm{q}}(m,n)}I$. Hence, a Bell violation occurs whenever there is a $\mathsf{p}\in C_{\mathrm{q}}(m,n)$ for which $I>\eta^{\mathrm{L}}$. A violation of a Bell inequality by non-communicating provers employing a quantum model is an indication of entanglement between provers.
@@
Typically if $\eta^{\mathrm{L}}$ is known for a given $I$, the main question is finding an upper bound on $\eta^{\mathrm{Q}}$. In this case, one often considers the Bell operator [^3] For a more mathematically rigorous treatment of such Bell operators and the SOS approach consult [CMM+24]. $S=\sum_{abxy}w_{abxy}\,M_{a|x}N_{b|y}$, and $\langle S\rangle=\langle\Psi|S|\Psi\rangle$ its quantum expectation with respect to $|\Psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Remark that since bipartite models with separable quantum states generate the classical correlations $C_{\mathrm{q}}(m,n)$, $\langle\Psi|S|\Psi\rangle\leq\eta^{\mathrm{L}}$ whenever $|\Psi\rangle$ is separable (unentangled). However, it’s possible that there could be entangled states for which $\langle\Psi^{\prime}|S|\Psi^{\prime}\rangle>\eta^{\mathrm{L}}$. Hence, given a Bell operator $S$, we can recover the maximum classical and quantum values $\eta^{\mathrm{L}}=\sup_{\mathrm{C}\in\mathcal{Q}(n,m)}\langle S\rangle$ and $\eta^{\mathrm{Q}}=\sup_{\mathrm{Q}\in\mathcal{Q}_{n,m}}\langle S\rangle$ respectively. Technically, we have not fixed the dimensions of the Bell operator as we want to consider any finite-dimensional model. Hence, the $\sup$ is implicitly also optimizing over all finite-dimensional Hilbert spaces $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$.
An approach to establishing upper bounds on $\langle S\rangle$ is using sum-of-squares techniques. Let $S$ be a Bell operator and $\eta^{\prime}>0$. The shifted Bell operator $\eta^{\prime}\mathbbm{1}-S$ admits a sum-of-squares (SOS) decomposition if there exists a set of complex polynomials $\{P_{i}\}_{i\in\mathcal{I}}$ in the elements $\{M_{a|x},N_{b|y}:a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$ satisfying
$\eta^{\prime}\mathbbm{1}-S=\sum_{i\in\mathcal{I}}P_{i}^{\dagger}P_{i}.$ (2.5)
@@
The existence of an SOS decomposition for the operator $\eta^{\prime}\mathbbm{1}-S$ implies that $\eta^{\prime}\mathbbm{1}-S$ is positive, and therefore $\eta^{\prime}$ is an upper bound on the maximum quantum value of $\langle S\rangle$. Additionally, if $\eta^{\prime}$ is achievable by a bipartite model, then we write $\eta^{\prime}=\eta^{\mathrm{Q}}$. In this case, the shifted Bell operator is $\bar{S}=\eta^{\mathrm{Q}}\mathbbm{1}-S$, and observing $\langle\Psi|\bar{S}|\Psi\rangle=0$ implies the constraints $P_{i}|\Psi\rangle=0$ for all $i\in\mathcal{I}$; these constraints can often be used to infer the algebraic structure (rigidity) of the measurements $\{M_{a|x}\},\{N_{b|y}\}$ which achieve $\langle S\rangle=\eta^{\mathrm{Q}}$.
3 Compiled Bell inequalities
@@
The compilation procedure of a Bell scenario is essentially the same as the procedure for compiling nonlocal games outlined in [KLVY23]. Let $\mathcal{S}=(\mathcal{X},\mathcal{Y},\mathcal{A},\mathcal{B},\pi)$ be a 2-prover Bell scenario and $\mathsf{Q}$ a quantum homomorphic encryption scheme with security parameter $\lambda$. A compiled Bell scenario is the following 2-round single-prover scenario. To setup, the verifier samples a pair of inputs $(x,y)\in\mathcal{X}\times\mathcal{Y}$ according to the distribution $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$. Additionally, the verifier samples a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ and encrypts the first input as the ciphertext $\chi\leftarrow\mathsf{Enc}(\mathsf{sk},x)$.
@@
In the compiled scenario, the prover acts as a single Hilbert space $\widetilde{\mathcal{H}}$ and prepares an initial state $|\Psi\rangle\in\widetilde{\mathcal{H}}$. Given a secret key $\mathsf{sk}$, the first round of the protocol is characterized by a family of POVMs $\{\{\widetilde{M}_{\alpha|\chi}\}_{\alpha\in\bar{\mathcal{A}}}\}_{\chi\in\bar{\mathcal{X}}}$ and unitaries $\{U_{\alpha,\chi}\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}$, where $\bar{\mathcal{X}}$ and $\bar{\mathcal{A}}$ are the set of all ciphertexts of the first round input and output, respectively. Unlike in the bipartite setting, we must account for unitary operations applied to the post-measurement state in the first round. With this in mind, we denote the sub-normalized post-measurement state given the measurement over ciphertext $\chi$ and encrypted outcome $\alpha$ by
$U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}|\Psi\rangle\coloneqq|\Psi_{\alpha|\chi}\rangle.$ (3.1)
@@
Note that these vectors are sub-normalized. In particular, the probability of obtaining $\alpha\in\bar{\mathcal{A}}$ given $\chi\in\bar{\mathcal{X}}$ is given by $\langle\Psi_{\alpha|\chi}|\Psi_{\alpha|\chi}\rangle$. In the second round, the device makes a POVM measurement $\{\{N_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}$, where the resulting conditional probability is given by
$\langle\Psi|\widetilde{M}_{\alpha|\chi}^{\dagger}U_{\alpha,\chi}^{\dagger}N_{b|y}U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}|\Psi\rangle=\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle,$ (3.2)
@@
for a fixed $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$, ciphertexts $\chi\in\bar{\mathcal{X}}$, $\alpha\in\bar{\mathcal{A}}$, and plaintexts $y\in\mathcal{Y}$, $b\in\mathcal{B}$.
@@
To summarize, for a fixed QHE scheme, a compiled (quantum) model is given by a tuple
$\widetilde{\mathrm{Q}}=(\widetilde{\mathcal{H}},\{|\Psi_{\alpha|\chi}\rangle\},\{\{N_{b|y}\}),$ (3.3)
@@
where all the relevant measurements and states are obtained by a QPT procedure. We remark that one can consider a description of the model which includes the initial state $|\Psi\rangle$ and the operators $\{U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}\}$, rather than the post-measurement states $|\Psi_{\alpha|\chi}\rangle$. Hence, $\widetilde{\mathrm{Q}}$ is really a coarser description of a quantum model in the compiled setting. The joint distribution of the outcomes after both rounds is given by
@@
$p(a,b,|x,y)=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle.$ (3.4)
Throughout the work, we will often suppress the expectation over secret keys $\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}$ and simply write the expectation for a fixed key.

### v3

@@ 2.1 Mathematical notation
Throughout the article, Hilbert spaces are denoted by $\mathcal{H}$, and are assumed to be finite-dimensional unless explicitly stated otherwise. Elements of $\mathcal{H}$ are denoted by $|v\rangle\in\mathcal{H}$, where the inner product $\langle u|v\rangle$ for $|v\rangle,|u\rangle\in\mathcal{H}$ is linear in the second argument and defines the vector norm $\||v\rangle\|=\sqrt{\langle v|v\rangle}$. Quantum pure states are the norm 1 elements of $\mathcal{H}$. In this work, $\mathbb{B}(\mathcal{H})$ denotes the unital $\dagger$-algebra of bounded linear operators on $\mathcal{H}$ with norm $\|M\|^{2}_{\mathrm{op}}=\sup_{|v\rangle\in\mathcal{H},|v\rangle\neq 0}\langle v|M^{\dagger}M|v\rangle/\langle v|v\rangle$. We also write $\|A\|_{2}=\sqrt{{\mathrm{t}r}(A^{\dagger}A)}$ to denote the Schatten 2-norm for $A\in\mathbb{B}(\mathbb{C}^{d})\cong M_{d}(\mathbb{C})$. The unit in $\mathbb{B}(\mathcal{H})$ is denoted by $\mathbbm{1}$, and we write $|M|=\sqrt{M^{\dagger}M}$ for the positive part of $M\in\mathbb{B}(\mathcal{H})$. Given a finite set $\mathcal{A}$, a collection of positive operators $\{M_{a}\geq 0:a\in\mathcal{A}\}$ with the property that $\sum_{a\in\mathcal{A}}M_{a}=\mathbbm{1}$, is called a POVM over $\mathcal{A}$. When the operators in a POVM are orthogonal projections, we call it a PVM. Given a random variable $X$, which takes values $X=x\in\mathcal{X}$ according to a distribution $\mu:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\sum_{x\in X}\mu(x)=1$, we denote the expectation of $X$ by $\mathop{\mathbb{E}}[X]=\sum_{x\in\mathcal{X}}\mu(x)\cdot x$. For $a,b\in\mathbb{R}$ and $\delta>0$, $a\approx_{\delta}b$ is short for $|a-b|\leq\delta$. A function $\text{negl}:\mathbb{N}\rightarrow\mathbb{R}$ is called negligible if for all $k\in\mathbb{N}$ there exists $N\in\mathbb{N}$ such that for every $n\geq N$ it holds that $\text{negl}(n)\leq\frac{1}{n^{k}}$.
@@ 2.2 Bell scenarios, inequalities, and violations
Before we discuss compiled Bell inequalities, let us recall the bipartite case. Here we let $\mathcal{A},\mathcal{B},\mathcal{X},$ and $\mathcal{Y}$ be finite sets, with $|\mathcal{A}|=m_{A}$, $|\mathcal{B}|=m_{B}$, $|\mathcal{X}|=n_{A}$, and $|\mathcal{Y}|=n_{B}$. A bipartite Bell scenario is described by the tuple $\mathcal{S}=(\mathcal{A},\mathcal{B},\mathcal{X},\mathcal{Y},\pi)$, where $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a distribution over the measurement settings. In a scenario, each party receives an input $x\in\mathcal{X}$ (resp. $y\in\mathcal{Y}$) sampled according to $\pi$, and returns outputs $a\in\mathcal{A}$ (resp. $b\in\mathcal{B}$). The parties are non-communicating, and therefore cannot coordinate their outputs. The behaviour of the provers is characterized by a correlation, a set of conditional probabilities $\mathsf{p}=\{p(a,b|x,y):a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$, which is realized by an underlying physical theory or model. In the quantum setting, we allow the provers to share a bipartite quantum state, and say the correlation $\mathsf{p}$ is realized by a bipartite (quantum) model
@@
where $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ are Hilbert spaces, $\{M_{a|x}\}_{a\in\mathcal{A}}$ and $\{N_{b|y}\}_{b\in\mathcal{B}}$ are POVMs on $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ respectively, and $|\Psi\rangle_{AB}$ is a vector state in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. More generally, a correlation $\mathsf{p}$ is quantum (or an element of $C_{\mathrm{q}}(n_{A},n_{B},m_{A},m_{B})$) if there exists a bipartite model $\mathrm{Q}$ for which $\mathsf{p}$ can be realized via the Born rule as $p(a,b|x,y)=\langle\Psi|M_{a|x}\otimes N_{b|y}|\Psi\rangle$. We denote the class of bipartite (quantum) models by $\mathcal{Q}(n_{A},n_{B},m_{A},m_{B})$. From now on we will refer to such models simply as bipartite models.
In contrast to the set of quantum correlations, we have the collection of local correlations $C_{\mathrm{loc}}(n_{A},n_{B},m_{A},m_{B})$. These are the correlations $\{p(a,b|x,y)\}$ for which there exists a classical model, that is a probability distribution $\mu_{k}$ and a local distributions $p_{k}^{A}(a|x)$ and $p_{k}^{B}(b|y)$ such that $p(a,b|x,y)=\sum_{k}\mu_{k}\,p_{k}^{A}(a|x)\,p_{k}^{B}(b|y)$. We let $\mathrm{C}=(\mu_{k},\{p_{k}^{A}\},\{p_{k}^{B}\})$ denote a classical model and let $\mathcal{L}(n_{A},n_{B},m_{A},m_{B})$ denote the class of all classical models. In what follows we consider Bell scenarios where $n_{A}=n_{B}=n$, and $m_{A}=m_{B}=m$. With this notation Bell’s theorem [Bel64] states that $C_{\text{loc}}(2,2)$ is a strict subset of $C_{\text{q}}(2,2)$.
@@
$I=\sum_{a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}}w_{abxy}\,p(a,b|x,y),$ (2.2)
@@
for coefficients $w_{abxy}\in\mathbb{R}$. A Bell inequality is a functional $I$ and a bound $\eta>0$ such that $I\leq\eta$ for all $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. Given a functional $I$, the classical value is the maximal value achieved by the classical correlations $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. We denote this value by $\eta^{\mathrm{L}}:=\sup_{\mathsf{p}\in C_{\mathrm{loc}}(m,n)}I$. The quantum value for $I$ is the maximal value achieved by the set of quantum correlations $\mathsf{p}\in C_{\mathrm{q}}(m,n)$, and we denote the quantum value on $I$ by $\eta^{\mathrm{Q}}:=\sup_{\mathsf{p}\in C_{\mathrm{q}}(m,n)}I$. Hence, a Bell violation occurs whenever there is a $\mathsf{p}\in C_{\mathrm{q}}(m,n)$ for which $I>\eta^{\mathrm{L}}$. A violation of a Bell inequality by non-communicating provers employing a quantum model is an indication of entanglement between provers.
@@
Typically when $\eta^{\mathrm{L}}$ is known for a given $I$, the main challenge is finding an upper bound on $\eta^{\mathrm{Q}}$. In this case, one often considers the Bell operator [^1] For a more mathematically rigorous treatment of Bell operators and the SOS approach consult [CMM+24]. $S=\sum_{abxy}w_{abxy}\,M_{a|x}\otimes N_{b|y}$, and $\langle S\rangle=\langle\Psi|S|\Psi\rangle$ its quantum expectation with respect to $|\Psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Since bipartite models with separable quantum states generate the classical correlations $C_{\mathrm{q}}(m,n)$, $\langle\Psi|S|\Psi\rangle\leq\eta^{\mathrm{L}}$ whenever $|\Psi\rangle$ is separable (unentangled). However, it’s possible that there could be entangled states for which $\langle\Psi^{\prime}|S|\Psi^{\prime}\rangle>\eta^{\mathrm{L}}$. Hence, given a Bell operator $S$, we can recover the maximum classical and quantum values $\eta^{\mathrm{L}}=\sup_{\mathrm{C}\in\mathcal{L}(n,m)}\langle S\rangle$ and $\eta^{\mathrm{Q}}=\sup_{\mathrm{Q}\in\mathcal{Q}(n,m)}\langle S\rangle$ respectively. Technically, we have not fixed the dimensions of the Bell operator as we want to consider any finite-dimensional model. Hence, the supremum is implicitly over all finite-dimensional Hilbert spaces $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$.
@@
An approach to establishing upper bounds on $\langle S\rangle$ is using sum-of-squares techniques. Let $S$ be a Bell operator and $\eta^{\prime}>0$. The shifted Bell operator $\eta^{\prime}\mathbbm{1}-S$ admits a sum-of-squares (SOS) decomposition if there exists a set of polynomials $\{P_{i}\}_{i\in\mathcal{I}}$ in the elements $\{M_{a|x},N_{b|y}:a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$ satisfying $\eta^{\prime}\mathbbm{1}-S=\sum_{i\in\mathcal{I}}P_{i}^{\dagger}P_{i}$. The existence of an SOS decomposition for the operator $\eta^{\prime}\mathbbm{1}-S$ implies that $\eta^{\prime}\mathbbm{1}-S$ is positive, and therefore $\eta^{\prime}$ is an upper bound on the maximum quantum value of $\langle S\rangle$. Additionally, if $\eta^{\prime}$ is achievable by a bipartite model, then we write $\eta^{\prime}=\eta^{\mathrm{Q}}$. In this case, the shifted Bell operator is $\bar{S}=\eta^{\mathrm{Q}}\mathbbm{1}-S$, and observing $\langle\Psi|\bar{S}|\Psi\rangle=0$ implies the constraints $P_{i}|\Psi\rangle=0$ for all $i\in\mathcal{I}$; these constraints can often be used to infer the algebraic structure (rigidity) of the measurements $\{M_{a|x}\}_{a\in\mathcal{A},x\in\mathcal{X}},\{N_{b|y}\}_{b\in\mathcal{B},y\in\mathcal{Y}}$ which achieve $\langle S\rangle=\eta^{\mathrm{Q}}$.
3 Compiled Bell scenarios
The compilation procedure of a Bell scenario is essentially the same as the procedure for compiling nonlocal games outlined in [KLVY23]. Let $\mathcal{S}=(\mathcal{X},\mathcal{Y},\mathcal{A},\mathcal{B},\pi)$ be a 2-prover Bell scenario and fix a quantum homomorphic encryption scheme with security against quantum distinguishers and correctness with respect to auxiliary input. Readers unfamiliar with QHE schemes and these properties can refer to Definition A.2 found in the appendix.
@@ 3 Compiled Bell scenarios
A compiled Bell scenario is the following 2-round single-prover scenario. To setup, the verifier samples a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$. Then, the verifier samples a pair of inputs $(x,y)\in\mathcal{X}\times\mathcal{Y}$ according to the distribution $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$, and encrypts the first input as the ciphertext $\chi\leftarrow\mathsf{Enc}(\mathsf{sk},x)$.
@@
In the compiled scenario, for a chosen security parameter $\lambda$, the prover prepares an initial quantum polynomial time (QPT) preparable state $|\Psi^{(\lambda)}\rangle\in\widetilde{\mathcal{H}}^{(\lambda)}$ where $\widetilde{\mathcal{H}}^{(\lambda)}$ is a single Hilbert space (see Definition A.1 for details on efficient quantum procedures). Then, the first round of the protocol is characterized by a family of POVMs $\{\{\widetilde{M}_{\alpha|\chi}^{(\lambda)}\}_{\alpha\in\bar{\mathcal{A}}}\}_{\chi\in\bar{\mathcal{X}}}$ and unitaries $\{U_{\alpha,\chi}^{(\lambda)}\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}$, where $\bar{\mathcal{X}}$ and $\bar{\mathcal{A}}$ are the set of all valid ciphertexts of the first round input and output, respectively. Unlike in the bipartite setting, we must account for unitary operations applied to the post-measurement state in the first round. With this in mind, we denote the sub-normalized post-measurement state given the measurement over ciphertext $\chi$ and encrypted outcome $\alpha$ by
$U_{\alpha,\chi}^{(\lambda)}\widetilde{M}_{\alpha|\chi}^{(\lambda)}|\Psi^{(\lambda)}\rangle=:|\Psi_{\alpha|\chi}^{(\lambda)}\rangle.$ (3.1)
@@
Note that these vectors are sub-normalized. In particular, the probability of obtaining $\alpha\in\bar{\mathcal{A}}$ given $\chi\in\bar{\mathcal{X}}$ is given by $\langle\Psi_{\alpha|\chi}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle$. In the second round, the device makes a POVM measurement $\{\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}$, where the resulting conditional probability is given by
$\langle\Psi^{(\lambda)}|{\widetilde{M}_{\alpha|\chi}^{(\lambda)\dagger}}{U_{\alpha,\chi}^{(\lambda)\dagger}}N_{b|y}^{(\lambda)}U_{\alpha,\chi}^{(\lambda)}\widetilde{M}_{\alpha|\chi}^{(\lambda)}|\Psi^{(\lambda)}\rangle=\langle\Psi_{\alpha|\chi}^{(\lambda)}|N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle,$ (3.2)
@@
for a fixed, $\lambda\in\mathbb{N}$, $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$, ciphertexts $\chi\in\bar{\mathcal{X}}$, $\alpha\in\bar{\mathcal{A}}$, and plaintexts $y\in\mathcal{Y}$, $b\in\mathcal{B}$.
@@
To summarize, for a fixed QHE scheme, $\lambda\in\mathbb{N}$, a compiled (quantum) model is given by a tuple
$\widetilde{\mathrm{Q}}^{(\lambda)}=(\widetilde{\mathcal{H}}^{(\lambda)},\{|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}},\{\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}),$ (3.3)
@@
where all the relevant measurements and states are obtained by some QPT procedure. We remark that one can consider a description of the model which includes the initial state $|\Psi^{(\lambda)}\rangle$ and the operators $\{U_{\alpha,\chi}^{(\lambda)}\widetilde{M}_{\alpha|\chi}^{(\lambda)}\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}$, rather than the post-measurement states $|\Psi_{\alpha|\chi}^{(\lambda)}\rangle$. Hence, $\widetilde{\mathrm{Q}}^{(\lambda)}$ is really a coarse description of a quantum model in the compiled setting. The joint distribution of the outcomes after both rounds is given by
@@
$p^{(\lambda)}(a,b|x,y)=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}^{(\lambda)}|N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle.$ (3.4)